Current Incidents


All systems are operating normally.


Past Incidents


Los Angeles Resolved

Today we have finished all the corrections, and the migration has finally been completed. From now on, we expect to maintain 100% uptime, which is our main objective in Los Angeles. We appreciate your patience and apologize for any inconvenience caused.

Incident reported: 24/11/2025 02:40 pm
Update History

Fixed

November 24, 2025 - 06:20 PM STATUS SET TO: Resolved

Service Degradation Due to Storage Issue Resolved

We are currently investigating a storage-related issue affecting the Dallas 1 node.
Our team is performing an urgent review to ensure data integrity and restore full functionality as quickly as possible.


Services hosted on this node may experience delays or temporary unavailability during the diagnostic process.
Further updates will be provided as soon as additional information becomes available.

Incident reported: 21/11/2025 11:58 am
Update History

Fixed

November 22, 2025 - 04:49 PM STATUS SET TO: Resolved

We have managed to restore the affected VPSs. For now, only these VMs are experiencing problems. We remind you to always back up your important data. We do not back up your VMs.

209.182.100.115
23.173.152.254
209.182.100.178

November 21, 2025 - 02:53 PM STATUS SET TO: Investigating

INCIDENT UPDATE - November 21, 2025

ISSUE IDENTIFIED:
Critical hardware failure detected on storage drive (nvme0n1) in Dallas datacenter.

AFFECTED SERVICES:
13 VPS instances with the following IP addresses:
- 23.173.152.249
- 23.173.152.43
- 23.173.152.247
- 23.173.152.79
- 23.173.152.236
- 23.173.152.160
- 209.182.100.205
- 23.173.152.7
- 209.182.100.110
- 209.182.100.20
- 209.182.100.233

CURRENT STATUS:
- All affected VPS instances are offline
- Data on failed drive is unrecoverable
- Infrastructure restoration in progress


NEXT STEPS:
Technical team is currently removing failed drive from volume group and stabilizing remaining infrastructure. Estimated completion: 2-3 hours.

November 21, 2025 - 12:15 PM STATUS SET TO: Investigating

Migration Notice LOS ANGELES – Phase 1 Resolved

We are beginning the first phase of our migration process. In this stage, five servers are being moved, and some Ontime services may experience temporary disruptions.


In the second phase, we will be relocating the main server that hosts the VPS services. We will inform you in advance of the exact date when this transfer will take place.


 


Thank you for your understanding during this process.

Incident reported: 12/11/2025 01:47 pm
Update History

Fixed

November 19, 2025 - 12:34 PM STATUS SET TO: Resolved

The VPS host server is now online, and all services are operational. Please verify that your VPS is functioning correctly. We are currently working on restoring the dedicated servers and the Ontime services server. Once all systems are restored, we will proceed to add service credits to all affected accounts.

November 19, 2025 - 09:58 AM STATUS SET TO: Investigating

We have already installed your servers in the rack, but we are currently experiencing an issue with the switch, which the datacenter is working to resolve.

November 19, 2025 - 07:36 AM STATUS SET TO: Investigating

We are on Sharktech time. This will take place today at 8:00 PM PDT.

November 18, 2025 - 11:34 AM STATUS SET TO: Investigating

Today, Sharktech already has our hardware at their facility, and we are now just waiting for it to be installed so we can come back online in Los Angeles. We hope to have no further issues.

November 17, 2025 - 06:19 PM STATUS SET TO: Investigating

Unfortunately, the rain continues in Los Angeles, making it impossible for LayerHost to move the servers to Sharktech. We hope to have better weather tomorrow or Wednesday. Please remember that all invoices will be paid automatically. We appreciate your understanding and patience during this process.

November 17, 2025 - 11:09 AM STATUS SET TO: Investigating

This is taking longer than expected. Due to heavy rain, the datacenter has not been able to move the servers to Sharktech. We still have not received a response from them confirming the new transfer date or the rack allocation for these servers

November 15, 2025 - 08:13 PM STATUS SET TO: Investigating

All services in Los Angeles are currently offline while the migration is in progress. This outage will remain in place for the duration of the relocation work.

We kindly ask for your patience during this process, as this migration is being carried out to improve service reliability and overall performance.

November 14, 2025 - 08:06 AM STATUS SET TO: Investigating

The first phase of the migration will take place tonight, during the night hours in Los Angeles, California. During this period, some services may experience temporary unavailability. We appreciate your understanding as we work to complete this transition as smoothly as possible.

November 13, 2025 - 12:27 PM STATUS SET TO: Investigating

Los Angeles Datacenter Outage Resolved

We are currently experiencing issues in Los Angeles. The data center is performing a review to restore services.

Incident reported: 12/11/2025 11:33 am
Update History

This has been resolved. If you have any issues, please open a support ticket.

November 12, 2025 - 01:26 PM STATUS SET TO: Resolved

DDOS attack DALLAS Resolved

As of 6pm EST on Sunday, October 26th, 2025, we continue to receive reports of short duration "blips" at all of our locations Dallas.   We are experiencing a "carpet bomb" style DDOS attack and continue to adjust our mitigation techniques to counter it.   We expect a full resolution shortly.

Incident reported: 26/10/2025 05:55 pm
Update History

this is considered resolved

October 28, 2025 - 09:55 AM STATUS SET TO: Resolved

Update 10/27 9:30AM EST:

After making more changes, we have seen the attacks fully subside for the longest stretch yet. We are cautiously optimistic that the issue is now resolved. We will keep the network status page open as we work through any remaining gremlins.

October 27, 2025 - 09:14 AM STATUS SET TO: Investigating

Update 11:44PM EST. We have successfully mitigated the majority of the attack in our Dallas and Ashburn POPs. We are continuing to fine tune those POPs as well as work on mitigation in our other locations. We appreciate your patience as we mitigate this complex attack.


Update 10/27 6:15AM EST:

We want to provide an update on the ongoing distributed denial of service (DDoS) attacks affecting our network. While we’ve made progress stabilizing some regions overnight, the attack remains active and continues to shift patterns. Our network and engineering teams have been working around the clock to keep services as stable as possible while preparing a long-term, comprehensive mitigation plan. We have recently learned that we are not the only target of this attack and other nearby providers are also battling it.

This morning, we are continuing with a plan for a full-scale migration of our prefixes to CosmicGuard and MagicTransit, two industry-leading DDoS protection platforms. This process is complex and will take time as we carefully transition each network segment to ensure stability and minimize further disruption.

We want to sincerely apologize for the ongoing interruptions and thank all of our customers for their continued patience and understanding. While we cannot provide an exact ETA for full restoration, this transition represents a major step toward a durable and permanent resolution.



We remain confident that once the migration is complete, our network will be significantly more resilient against this and future large-scale attacks. Further updates will be posted here as progress continues throughout the day.

October 27, 2025 - 07:45 AM STATUS SET TO: Investigating